OpenStudy (gabylovesyou):

If f(x) = x + 2, find f(3).

jimthompson5910 (jim_thompson5910):

Replace 'x' with 3 and evaluate f(x) = x + 2 f(3) = 3 + 2

OpenStudy (gabylovesyou):

f(3) = 5

jimthompson5910 (jim_thompson5910):

yes, you got it

OpenStudy (gabylovesyou):

thank you

jimthompson5910 (jim_thompson5910):

you're welcome